# {NoHyper}footnotetext: *Co-first author

### v1

@@ 3.4 Precise Numeric Reasoning in LLMs
Graphics programming requires the precise estimation of continuous quantities such as location and orientation, along with a comprehension of Euclidean space. This requirement extends beyond the coarse, semantic-level spatial reasoning (e.g. “left,” “in front of”) for which LLMs are typically employed, in tasks such as VQA (Antol et al., 2015). Estimating continuous values through character-based outputs essentially transforms the task into a discrete, combinatorial challenge. In this loss space, prediction errors do not reflect real metric distances – a ground truth value of ‘4’ is considered as close to a prediction of ‘3’ as it is to ‘8,’ highlighting the inherent limitations of this approach for tasks demanding high numerical precision.
@@
We systematically investigate the behavior of character-based and numeric IG-LLM variants for precise spatial reasoning in Sec. 4.2. Our empirical findings support our intuition regarding the limitations of the character-based output and demonstrate that the numeric head enables strong generalization when the testing samples are OOD in parameter space. These differences are highlighted throughout our evaluation.
@@ 4.1 Compositional Generalization on CLEVR
An extension to CLEVR, known as CLEVR-CoGenT (Johnson et al., 2017), serves as a benchmark for evaluating the compositional-generalization capabilities of VQA models. This benchmark assesses the model’s ability to answer questions about scenes containing objects with unseen combinations of attributes. During training, the dataset is structured such that particular types of objects are only assigned specific combinations of attributes (e.g. blue cubes and red cylinders), while the testing data includes objects with attribute combinations not seen during training (e.g. red cubes and blue cylinders). We adapt this VQA dataset to our inverse-graphics problem domain, employing it for three primary purposes: 1) demonstrating that LLMs can effectively perform inverse graphics by testing on in-distribution (ID) data; 2) illustrating that LLMs exhibit robust compositional generalization to OOD data, while the baseline approach in NS-VQA (Yi et al., 2018) struggles in this setting; and 3) exploring the data-efficiency of our framework.
@@
Evaluation Metrics. To evaluate attribute-recognition accuracy, we employ linear-sum assignment on pairwise Euclidean distances to match predicted and ground-truth objects. However, since attribute-recognition accuracy does not account for missing or duplicated objects, we also evaluate the method’s ability to produce accurate counts by computing the mean-absolute counting error between the predicted and ground-truth object sets across scenes (Count).
@@ 4.2 Numeric Parameter-Space Generalization
Figure 5: 2D Parameter-Space Generalization. (Sec. 4.2.1) (a) Training positions are sampled from the checkerboard. When evaluated on images with uniformly sampled positions, the char-based model fails to generalize outside the training distribution (b) while the float-based model effectively interpolates samples (c). Randomly-sampled testing locations are shown in red and the corresponding predictions in blue. (d) shows that while both methods well-estimate samples from the ID condition, the char-based model struggles to generalize. (e) shows a plot of the model’s validation MSE as a function of the number of training steps. We observe that the training of the float-based model is much smoother and converges quickly.
@@
We begin by scrutinizing the framework’s ability to generalize in 2D parameter space across range gaps. To accomplish this, we create a dataset comprising 10k images, each featuring a red dot on a white background. During training, the model is shown images where the location of the dot is sampled from a sparse checkerboard grid, as shown in Fig. 5(a). During evaluation, the model is shown 1k images where the dot’s location is uniformly sampled across the square; points lying outside the checkerboard are effectively OOD inputs.
Results. As shown in Fig. 5(b), the char-based model exhibits significant overfitting to the training distribution, consistently predicting dot locations restricted to the checkerboard distribution observed during training. In contrast, the float-based model is able to effectively generalize across parameter space, adapting to the uniformly sampled testing distribution during evaluation (Fig. 5(c)). Although the float-based model exhibits a slight positional bias toward predicting positions on the grid (as evidenced by the higher OOD error), the disparity in the ID–OOD validation-L2 performance gap of the char-based model is 14 times as high as that of the float-based model (Fig. 5(d)). Moreover, the validation MSE of the float-based model converges quickly to near zero, while the error of the char-based model is much less stable over time (Fig. 5(e)), suggesting that the float-based model learns smooth, low-dimensional representations of the space while the char-based variant may not.
@@ 4.2.2 SO(3) Parameter Space
Figure 6: SO(3) Range Gap. (Sec. 4.2.2) (a) Visualization of the SO(3) range-gap sampling space. For training, Euler rotation components are sampled from the blue regions. In OOD, samples are exclusively sampled from the red ranges. (b) The float-based model outperforms the char-based variant on validation data sampled from both the ID and OOD conditions.
@@
As depicted in Fig. 6(b), the error of the char-based model is 2.64 times higher than that of the float-based model when evaluated in-distribution. Upon testing in the OOD condition, the disparity is nearly consistent at 2.52 times that observed in the ID scenario, with the ID–OOD gap of the char-based model being 2.21 times that observed in the float-based variant. We attribute the superiority of the float-based model across both conditions to the increased data dimensionality. Additionally, the lesser performance decline observed when evaluating on the OOD training gaps further underscores the parameter-space efficiency of the float-based model.
@@
We examine the ability of our framework to scale in tackling a more challenging inverse-graphics task: that of 6-DoF-pose estimation. Our exploration begins with an evaluation on single-object images, encompassing both quantitative and qualitative assessments, where we illustrate the framework’s ability to generalize across visual domain shifts. We subsequently extend the setting to include more-complex (albeit, synthetic) multi-object scenes, demonstrating promising results for scene estimation, handling larger collections (>100) of diverse assets.
Table 2: Single-Object 6-DoF Results. (Sec. 4.3.1) When evaluating on ID data in the one-million-sample single-object 6-DoF eval, we observe little difference between models; both well-capture the distribution. Geod. represents geodesic distance in degrees.
@@ 4.3.1 Single-Object 6-DoF
Results. Tab. 2 illustrates that, under this non-data-constrained scenario, both model variants effectively capture the dynamics of the task The models both notably exhibit an order of magnitude lower positional error than in the CLEVR setting, despite the addition of 3D orientation and an additional positional dimension, and achieve rotational error 28% of that observed in the ID portion of the SO(3) range-gap evaluation. This reinforces the earlier observation from the CLEVR data-efficiency evaluation that, given sufficient data, the model variants exhibit a similar performance ceiling. Still, neither achieves the level of precision necessary to be directly constrained by the three-decimal-place discretization applied to numeric quantities throughout the evaluations nor the 16-bit training precision in the case of the float-based model. See Appendix A for further training details.
As part of our evaluation, we also qualitatively examine the ability of the model to transfer from the renders of toy planes with a solid-white background, on which it was fine-tuned, to estimating the pose and attributes of planes in real-world images. We provide qualitative samples of our model’s generalization to such images in Fig. 7. We observe encouraging generalization across the majority of images tested, despite the lack of augmentation or domain-specific inductive bias applied during the training process. However, it is difficult to quantitatively evaluate such model performance due to a lack of paired real-world data in line with our compositional task. As a proxy for such an evaluation, we introduce a synthetic setting in Sec. 4.3.2 to quantitatively evaluate the ability of our framework to generalize across visual domains.
@@ 4.3.2 Scene-Level 6-DoF
In this section, we explore the scalability of our framework to scene-level 6-DoF-pose estimation, featuring 3-5 objects per scene and a much-expanded array of assets. This experiment not only assesses performance under more-challenging conditions, but also enables a quantitative evaluation on the framework’s ability to generalize to scenes with OOD visual appearance.
Setting. We construct an expanded CLEVR-like image–scene dataset, incorporating objects sourced from ShapeNet (Chang et al., 2015). The dataset comprises 56 chair types, 35 sofa types, and 47 table types. We remove the size and material attributes used in CLEVR, but employ the expanded color set used in Sec. 4.3.1 to randomly color the objects. After doing so, the total number of possible combinations of attributes is 191-fold that used in the CLEVR-CoGenT experiment. Differing from previous evaluations, we also vary the pitch of the camera and the radius of its arc, but maintain a fixed camera focal point. Returning from the million-image single-object 6-DoF evaluation to a relatively data-constrained setting, we render 10k training images and evaluate the framework on three conditions, each with 1K images: (1) ID, which matches the training distribution of scenes with solid-colored objects; (2) OOD texture (OOD-T), where the same object assets are used as in ID but the objects are rendered with original ShapeNet textures instead of the randomly assigned solid colors; and (3) OOD encompassing both unseen objects and original ShapeNet textures (OOD-T+S). We use this to emulate the distribution shift of modeling real-world scenes, while facilitating quantitative evaluation.
Results. We observe that both approaches scale to the task, though the float-based model outperforms – or ties with – the char-based variant across evaluations (Tab. 3). This disparity is emphasized in the OOD-T+S setting where scene-level chamfer distance of the char-based model jumps from being approximately twice that of the float-based variant in the ID and OOD-T evaluations to being 5.67 times as much.
There is a decrease in performance observed when stepping to the OOD-T setting, which is most-strongly observed in the count error (x8 for both) and the shape-recognition accuracy (-20.49% in char and -15.05% in float). We empirically attribute this to the model occasionally explaining some multi-color textured objects using a composition of multiple, solid-color assets. Quantitatively supporting this, the performance decrease is not as strongly reflected in scene-level chamfer distance (x2.6 for both).
See Fig. 8 for samples reconstructions from OOD-T+S and Fig. S.8 for samples from the ID setting. We additionally test our model on real-world samples, but find that it fails to consistently generalize (Fig. S.1). We attribute this failure partially to limitations in the camera-position training distribution.
@@ 5 Discussion and Limitations
The primary limitation of our approach lies in that its expressiveness is constrained by the expressiveness of the training-data-generation framework. We demonstrated its ability to learn to compositionally disentangle images of scenes into constituent elements, reconstructing scenes under distribution shifts. However, reproducing scenes as text, it can reconstruct scenes containing unknown objects in OOD configurations, but it does so in terms of the objects – and language – it is trained with. If it does not know the name of asset chairs_0055, it will not be able to use it. Even if the model produces the name of a new color or shape from outside of the training data, the graphics engine rendering the LLM output must have an understanding of it in order to apply it.
In contrast, the generality of our approach, which doesn’t incorporate special task-specific inductive biases, allows it to scale with the diversity of the training data or the expressivity of the code format. Future work may explore more-scalable training-data generators or integrate self-supervision techniques to enable learning from unlabeled images. While we employ a relatively straightforward object-centric code representation across experiments for simplicity, more-expressive scene representations should also be explored.
Our evaluation scenes feature only minor object occlusions and are relatively simple. While a generic next-token objective paired with MSE float supervision sufficed for these scenarios, addressing harder-to-disentangle scenes may require a trade-off between generality and inductive bias, to incorporate additional supervision.
@@ 6 Conclusion
In this work, we investigated the ability of LLMs to solve inverse-graphics challenges. Introducing the Inverse-Graphics Large-Language-Model (IG-LLM) framework, we demonstrated that the broad generalization and reasoning capabilities of LLMs can be harnessed to facilitate inverse-graphics tasks. Through extensive evaluation, we assessed the model’s capacity to generalize out-of-domain, revealing its ability to abstract scene elements compositionally. We additionally explored the integration of a numeric head to adapt LLMs for continuous metric-value estimation, providing enhanced generalization and smoother training dynamics. Our quantitative analyses demonstrate its ability to generalize compositionally (Sec. 4.1), in parameter space (Sec. 4.2), and across visual domains (Sec. 4.3). Our investigation demonstrates the ability of IG-LLM to leverage the general knowledge of LLMs in solving inverse-graphics problems, opening a new avenue for research.
@@ Appendix A Further Training Details
We finetune the LLaMA 1-based Vicuna 1.3 model [^1] https://huggingface.co/lmsys/vicuna-7b-v1.3 with LoRA (Hu et al., 2022a). We use the HuggingFace Transformers and PEFT libraries, along with DeepSpeed ZeRO-2 (Rajbhandari et al., 2020). In all experiments, we use a lora_r of 128, a lora_alpha of 256, a LoRA learning rate of 2e-05, a linear projector learning rate of 2e-05, a numeric head learning rate of 2e-04, and a cosine learning-rate schedule. All models are trained with an effective batch size of 32 with bfloat16 mixed-precision training. Both the cross-entropy next-token-prediction and mean-square-error losses are given a weight of 1.
The models for the CLEVR and parameter-space generalization experiments are trained for 40k steps. The 6-DoF pose-estimation models are trained for 200k steps.
We use the frozen CLIP visual tokenizer from  [^2] https://huggingface.co/openai/clip-vit-large-patch14-336. This CLIP variant has an input size of 336x336 pixels. For the CLEVR evaluation, we render images at the original size of 480x320 to ensure compatibility with NS-VQA, but pad and resize them for use with our model. For the remaining evaluations we directly render images at a resolution of 336x336.

### v2

@@ 3.4 Precise Numeric Reasoning in LLMs
Graphics programming requires the precise estimation of continuous quantities such as location and orientation, along with a comprehension of Euclidean space. This requirement extends beyond the coarse, semantic-level spatial reasoning (e.g., “left,” “in front of”) for which LLMs are typically employed, in tasks such as VQA (Antol et al., 2015). Estimating continuous values through character-based outputs essentially transforms the task into a discrete, combinatorial challenge. In this loss space, prediction errors do not reflect real metric distances – a ground truth value of ‘4’ is considered as close to a prediction of ‘3’ as it is to ‘8,’ highlighting a potential limitation inherent in this approach for learning a continous understanding.
@@
We systematically investigate the behavior of character-based and numeric IG-LLM variants for precise spatial reasoning in Sec. 4.2. Our empirical findings support our intuition regarding the limitations of the character-based output and demonstrate that the numeric head enables greater generalization when the testing samples are OOD in parameter space. These differences are highlighted throughout our evaluation.
@@ 4.1 Compositional Generalization on CLEVR
An extension to CLEVR, known as CLEVR-CoGenT (Johnson et al., 2017), serves as a benchmark for evaluating the compositional-generalization capabilities of VQA models. This benchmark assesses the model’s ability to answer questions about scenes containing objects with unseen combinations of attributes. During training, the dataset is structured such that particular types of objects are only assigned specific combinations of attributes (e.g., blue cubes and red cylinders), while the testing data includes objects with attribute combinations not seen during training (e.g., red cubes and blue cylinders). We adapt this VQA dataset to our inverse-graphics problem domain, employing it for three primary purposes: 1) demonstrating that LLMs can effectively perform inverse graphics by testing on in-distribution (ID) data; 2) illustrating that LLMs exhibit robust compositional generalization to OOD data, while the baseline approach in NS-VQA (Yi et al., 2018) struggles in this setting; and 3) exploring the data-efficiency of our framework.
@@
Evaluation Metrics. To evaluate attribute-recognition accuracy, we employ linear sum assignment on pairwise Euclidean distances to match predicted and ground-truth objects. However, since attribute-recognition accuracy does not account for missing or duplicated objects, we also evaluate the method’s ability to produce accurate counts by computing the mean-absolute counting error between the predicted and ground-truth object sets across scenes (Count). L2 represents the positional L2 norm and Size, Color, Mat., and Shape are percentages of attribute-recognition accuracy.
@@ 4.2 Numeric Parameter-Space Generalization
Figure 5: 2D Parameter-Space Generalization. (Sec. 4.2.1) (a) Training positions are sampled from the checkerboard. When evaluated on images with uniformly sampled positions, the char-based model fails to generalize outside the training distribution (b) while the float-based model effectively interpolates samples (c). Randomly sampled testing locations are shown in red and the corresponding predictions in blue. (d) shows that, while both methods well estimate samples from the ID condition, the char-based model struggles to generalize. (e) shows a plot of the model’s validation MSE as a function of the number of training steps. We observe that the training of the float-based model is much smoother and converges quickly.
@@
We begin by testing the framework’s ability to generalize in 2D parameter space across range gaps. To accomplish this, we create a dataset comprising 10k images, each featuring a red dot on a white background. During training, the model is shown images where the location of the dot is sampled from a sparse checkerboard grid, as shown in Fig. 5(a). During evaluation, the model is shown 1k images where the dot’s location is uniformly sampled across the square; points lying outside the checkerboard are effectively OOD inputs.
Results. As shown in Fig. 5(b), the char-based model exhibits strong overfitting to the training distribution, consistently predicting dot locations restricted to the checkerboard distribution observed during training. In contrast, the float-based model is able to effectively generalize across parameter space, adapting to the uniformly sampled testing distribution during evaluation (Fig. 5(c)). Although the float-based model exhibits a slight positional bias toward predicting positions on the grid (as evidenced by the higher OOD error), the disparity in the ID–OOD validation-L2 performance gap of the char-based model is 14 times as high as that of the float-based model (Fig. 5(d)). Moreover, the validation MSE of the float-based model converges quickly to near zero, while the error of the char-based model is much less stable over time (Fig. 5(e)), suggesting that the float-based model learns smooth, low-dimensional representations of the space while the char-based variant may not.
@@ 4.2.2 SO(3) Parameter Space
Figure 6: SO(3) Range Gap. (Sec. 4.2.2) (a) Visualization of the SO(3) range-gap sampling space. For training, Euler rotations are sampled from the blue regions. In OOD, rotations are exclusively sampled from the red ranges. (b) The float-based model outperforms the char-based variant on validation data sampled from both the ID and OOD conditions. Geod. represents geodesic distance in degrees.
@@
As depicted in Fig. 6(b), the error of the char-based model is 2.64 times higher than that of the float-based model when evaluated in-distribution. Upon testing in the OOD condition, the disparity is nearly consistent at 2.52 times that observed in the ID scenario, with the ID–OOD gap of the char-based model being 2.21 times that observed in the float-based variant. We attribute the superiority of the float-based model across both conditions to the increased data dimensionality. Additionally, the lesser performance decline observed when evaluating on the OOD gaps further underscores the parameter-space efficiency of the float-based model.
@@
We examine the ability of our framework to scale in tackling a more challenging inverse-graphics task: that of 6-DoF pose estimation. Our exploration begins with an evaluation on single-object images, encompassing both quantitative and qualitative assessments, where we illustrate the framework’s ability to generalize across visual domain shifts. We subsequently extend the setting to include more-complex (albeit, synthetic) multi-object scenes, demonstrating promising results for scene estimation, handling larger collections (>100) of diverse assets.
@@
4.3.1 Single-Object 6-DoF
Table 2: Single-Object 6-DoF Results. (Sec. 4.3.1) When evaluating on ID data in the one-million-sample single-object 6-DoF eval, we observe little difference between models; both well capture the distribution.
@@ 4.3.1 Single-Object 6-DoF
Results. Tab. 2 illustrates that, under this non-data-constrained scenario, both model variants effectively capture the dynamics of the task. The models both notably exhibit a positional error that is an order of magnitude lower than in the CLEVR setting, despite the addition of 3D orientation and an additional positional dimension. They also achieve rotational error that is 28% of that observed in the ID portion of the SO(3) range-gap evaluation. This reinforces the earlier observation from the CLEVR data-efficiency evaluation that, given sufficient data, the model variants exhibit a similar performance ceiling. Still, neither achieves the level of precision necessary to be directly constrained by the three-decimal-place discretization applied to numeric quantities throughout the evaluations nor the 16-bit training precision in the case of the float-based model. See Appendix A for further training details.
As part of our evaluation, we also qualitatively examine the ability of the model to transfer from the renders of toy planes with a solid-white background, on which it was fine-tuned, to estimating the pose and attributes of planes in real-world images. We provide qualitative samples of our model’s generalization to such images in Fig. 7. We observe encouraging generalization across the majority of images tested, despite the lack of augmentation or domain-specific inductive bias applied during the training process. However, it is difficult to quantitatively evaluate model performance due to a lack of paired real-world data in line with our compositional task. As a proxy for such an evaluation, we introduce a synthetic setting in Sec. 4.3.2 to quantitatively evaluate the ability of our framework to generalize across visual domains.
@@ 4.3.2 Scene-Level 6-DoF
In this section, we explore the scalability of our framework to scene-level 6-DoF-pose estimation, featuring 3-5 objects per scene and a much-expanded array of assets. This experiment not only assesses performance under more-challenging conditions but also enables a quantitative evaluation on the framework’s ability to generalize to scenes with OOD visual appearances.
Setting. We construct an expanded CLEVR-like image–scene dataset, incorporating objects sourced from ShapeNet (Chang et al., 2015). The dataset comprises 56 chair types, 35 sofa types, and 47 table types. We remove the size and material attributes used in CLEVR but employ the expanded color set used in Sec. 4.3.1 to randomly color the objects. After doing so, the total number of possible combinations of attributes is 191-fold that used in the CLEVR-CoGenT experiment. Differing from previous evaluations, we also vary the pitch of the camera and the radius of its arc but maintain a fixed camera focal point. Returning from the million-image single-object 6-DoF evaluation, we render 100k training images and evaluate the framework on three conditions, each with 1K images: (1) ID, which matches the training distribution of scenes with solid-colored objects; (2) OOD texture (OOD-T), where the same object assets are used as in ID but the objects are rendered with original ShapeNet textures instead of randomly assigned solid colors; and (3) OOD encompassing both unseen objects and original ShapeNet textures (OOD-T+S). We use this to emulate the distribution shift of modeling real-world scenes while facilitating quantitative evaluation.
Results. We observe that both approaches scale to the task, though the float-based model outperforms – or ties with – the char-based variant across evaluations (Tab. 3).
There is a decrease in performance observed when stepping to the OOD-T setting, which is most-strongly observed in the count error (x4–5) and the shape-recognition accuracy (-15.07% in char and -12.38% in float). We empirically attribute this to the model occasionally explaining some multi-color textured objects using a composition of multiple, solid-color assets. Quantitatively supporting this, the performance decrease is not as strongly reflected in scene-level chamfer distance (x1.8–2.1).
See Fig. 8 for sample reconstructions from OOD-T+S and Fig. S.8 for samples from the ID setting. We additionally test our model on real-world samples but find that it fails to consistently generalize (Fig. S.1). We attribute this failure as likely due primarily to limitations in the training camera-position distribution.
@@ 5 Discussion and Limitations
The primary limitation of our approach lies in that its expressiveness is constrained by the expressiveness of the training-data-generation framework. We demonstrated its ability to learn to compositionally disentangle images of scenes into constituent elements, reconstructing scenes under distribution shifts. However, reproducing scenes as text, it can reconstruct scenes containing unknown objects in OOD configurations, but it does so in terms of the objects – and language – it is trained with. If it does not know the name of asset chairs_0055, it will not be able to use it. Even if the model produces the name of a new color or shape from outside the training data, the graphics engine rendering the LLM output must have an understanding of it in order to apply it.
In contrast, the generality of our approach, which doesn’t incorporate special task-specific inductive biases, allows it to scale with the diversity of the training data or the expressivity of the code format, beyond simple shapes. Future work may explore more-scalable training-data generators or integrate self-supervision techniques to enable learning from unlabeled images. In order to better focus our investigation on generalization across distributional shifts, we employ a relatively straightforward object-centric code representation, though more-expressive scene representations should also be explored.
Our evaluation scenes feature only minor object occlusions and are relatively simple. While a generic next-token-prediction objective paired with MSE float supervision sufficed for these scenarios, addressing harder-to-disentangle scenes may require a trade-off between generality and inductive bias to incorporate additional supervision such as differentiable rendering.
@@ 6 Conclusion
In this work, we investigated the ability of LLMs to solve inverse-graphics challenges. Introducing the Inverse-Graphics Large Language Model (IG-LLM) framework, we demonstrated that the broad generalization and reasoning capabilities of LLMs can be harnessed to facilitate inverse-graphics tasks. Through extensive evaluation, we assessed the model’s capacity to generalize out-of-domain, revealing its ability to abstract scene elements compositionally. We additionally explored the integration of a numeric head to adapt LLMs for continuous metric-value estimation, providing enhanced generalization and smoother training dynamics. Our quantitative analyses demonstrate its ability to generalize compositionally (Sec. 4.1), in parameter space (Sec. 4.2), and across visual domains (Sec. 4.3). Our investigation demonstrates the ability of IG-LLM to leverage the general knowledge of LLMs in solving inverse-graphics problems, opening a new avenue for research.
@@ Appendix A Further Training Details
We finetune the LLaMA 1-based Vicuna 1.3 model [^1] https://huggingface.co/lmsys/vicuna-7b-v1.3 with LoRA (Hu et al., 2022a). We use the HuggingFace Transformers and PEFT libraries, along with DeepSpeed ZeRO-2 (Rajbhandari et al., 2020). In all experiments, we use a lora_r of 128, a lora_alpha of 256, a LoRA learning rate of 2e-05, a linear projector learning rate of 2e-05, a numeric head learning rate of 2e-04, and a cosine learning-rate schedule. All models are trained with an effective batch size of 32 with bfloat16 mixed-precision training. Both the cross-entropy next-token-prediction and mean-square-error (MSE) losses are given a weight of 1.
The models for the CLEVR and parameter-space generalization experiments are trained for 40k steps. The single-object 6-DoF pose-estimation model is trained for 200k and the scene-level ShapeNet model for 500k steps.
We use the frozen CLIP visual tokenizer from  [^2] https://huggingface.co/openai/clip-vit-large-patch14-336. This CLIP variant has an input size of 336x336 pixels. For the CLEVR evaluation, we render images at the original size of 480x320 to ensure compatibility with NS-VQA, but we pad and resize them for use with our model. For the remaining evaluations, we directly render images at a resolution of 336x336.